# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning

### v1

@@ Abstract
Recent vision-language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to the over prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address this issue, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under one-word attacks, MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-MU. WARNING: There exist AI generations that may be offensive in nature.
@@ 1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images [1, 2, 3, 4, 5]. A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module [6, 7, 8, 9, 10]. While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities [11, 12, 13, 14, 15]. Specifically, images can function as a “foreign language” [16], creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety [16, 17, 18]. The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), provided that a high-quality, dual-modality curated safety fine-tuning dataset is available [14, 19, 20, 21, 22]. One compelling piece of evidence is that fine-tuning on VLGuard [23], a widely used VLM safety dataset, substantially improves robustness against unsafe queries and jailbreaking attacks. As demonstrated in [23], this enhancement surpasses the results obtained by fine-tuning on a “clean” dataset which removes the unsafe data. The surprising effectiveness of SFT on VLGuard has sparked growing interest in re-evaluating its reliability. Recent studies [18, 11, 24] have identified a downside of such safety fine-tuning, known as the over-prudence problem. This issue refers to the over-conservatism of VLMs after safety fine-tuning, where they unnecessarily reject responses when presented with benign queries.
However, the over-prudence suggests that these models may be overly safe, to avoid responses to any skeptical queries. Alternatively, the observed safety may be illusory, as current safety fine-tuning fails to ensure reliability, giving a false sense of safety. Thus, we ask:
@@
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed, fine-tuned on VLG uard [23]) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in the original unsafe query) can bypass the safety mechanism established through fine-tuning on VLGuard , even though the safeguarded VLM correctly rejects the original unsafe query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations given by fine-tuning dataset biases: Certain words become disproportionately associated with specific safety labels. For example, “Share” is strongly correlated with rejection responses, while “What” is highly associated with non-rejection responses. (d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO [25] and RMU [26] enhance robustness against attacks while reducing over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed [23] (“Prior work”).
@@
In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs. We uncover a “safety mirage” in VLM safety fine-tuning, where the seemingly robust safety performance after fine-tuning is primarily driven by spurious correlations between certain non-essential textual words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset. If an adversary identifies these spurious correlations, a simple one-word modification–which we refer to as the “one-word attack”–can effectively jailbreak safety fine-tuned VLMs, enabling them to regenerate unsafe content. Additionally, the input-rejection label shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs. Similar to the one-word attack, a one-word modification in text queries can readily activate the input-rejection shortcut at test time, causing the model to overgeneralize rejection responses and refuse to generate outputs even for benign queries. In Fig. 1(a)-(c), we provide a schematic overview illustrating: (a) The one-word attack on a safety fine-tuned VLM LLaVA-v1.5-7B-Mixed [23]; (b) The over-prudence issue by one-word modification; (c) The spurious correlations observed in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
@@
Building on our identification of spurious correlations in VLM safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU ensures that essential knowledge remains intact while avoiding unintended disruptions to causally unrelated information [27, 28, 29]. We propose adapting MU to VLM safety fine-tuning as a more robust alternative to traditional supervised approaches. Rather than enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) showcases the effectiveness of applying two LLM unlearning approaches, NPO [25] and RMU [26] in enhanced robustness against jailbreaking attacks and reduced over-prudence rates.
@@
① We revisit the problem of supervised safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases—specifically, spurious correlations between non-essential textual features and safety labels in the fine-tuning dataset.
② From an attack perspective, we show that safety fine-tuned VLMs are still susceptible to jailbreaking when adversaries exploit spurious correlations embedded in the fine-tuning dataset. We propose a simple and effective one-word attack by substituting highly frequent querying words associated with rejection responses with those linked to normal model outputs. Additionally, we show that these spurious correlations also contribute to over-prudence, causing fine-tuned VLMs to unnecessarily reject benign inputs.
③ From a defense perspective, we show that MU offers a promising solution to alleviate the effects of spurious correlations in fine-tuning data for VLM safety fine-tuning. The key rationale is that unlearning removes the influence of unsafe responses without relying on the spurious feature-label correlation present in the fine-tuning dataset.
④ We conduct extensive experiments across multiple VLM safety evaluation benchmarks, including VLGuard [23], SPA-VL [21], MM-SafetyBench [14], and FigStep [15], and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
@@ 2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs [6, 7, 8, 32, 33, 34, 1, 2, 4], safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies [35, 36, 37, 38], defense mechanisms [39, 40, 41], and robust evaluation datasets [42, 43, 44], VLMs introduce additional challenges due to the complexity of multimodal inputs [16, 17, 18], making them even more vulnerable to jailbreaking and adversarial manipulation [11, 12, 13, 14, 15]. Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs [15, 14]. Another strategy involves using gradient-based adversarial image generation [45, 46, 47, 48, 49] to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms [50, 51, 16, 52, 24]. The latter focuses on red-teaming dataset curation [14, 19, 20, 21, 22, 23, 53], enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
Machine unlearning in VLMs. MU [27, 28, 29, 54, 55] is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance [25, 26, 56, 57, 58, 59, 60]. In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms [61, 62, 63]. For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation [30, 31, 17]. Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
@@ 3 Preliminary and Problem Statement
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc [23, 64]). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreaking: The safety fine-tuned model originally produces rejection-based responses for unsafe queries; However, replacing the initial question word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by replacing “What” with “Share” can trigger unnecessary refusals even for benign queries.
@@
Previous works [16, 15, 14] highlight the need for safety alignment in VLMs to encompass both textual and imagery data. Consequently, many efforts [23, 21, 65] have focused on curating high-quality dual-modality safety datasets for VLMs. A notable benchmark dataset is VLGuard [23], which covers various text-image pairing scenarios, including unsafe cases where either the text is unsafe or both text and image are unsafe, as well as safe cases where both modalities are benign. Leveraging these curated safety datasets, recent works [23, 17, 18, 21] show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks [66, 67, 68]. In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
We begin by presenting the problem formulation of VLM safety fine-tuning. Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard). In addition, let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, which consists of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning. For a VLM parameterized by ${\bm{\theta}}$, the safety fine-tuning problem can be formulated as:
@@
where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and the regularization parameter $\gamma\geq 0$ strikes a balance between safety alignment and preserving performance on normal tasks.
Safety mirage: Motivation and problem of interest. Although the safety fine-tuned VLMs could overly reject queries even when presented with benign ones [18, 11, 24], the resulting safety performance against unsafe queries does not degrade; It remains highly robust even in the presence of some common jailbreaking attacks [23, 21].
The seemingly ‘robust’ safety performance observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement to Fig. 1, Fig. 2 provides additional motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. Consistent with Fig. 1, unsafe queries prefixed with innocuous question word “What” successfully bypass safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the replacement word ‘What’ or ‘Share’ is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
Examples in Figs. 1-2 suggest that fine-tuning VLMs on safety datasets may create a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries. Thus, our work focuses on the following key research questions: (a) What is the root cause of the “safety mirage” in VLM safety fine-tuning? (b) What can be improved to mitigate the “safety mirage”? We explore these questions in Sec. 4 and Sec. 5, respectively.

### v2

@@ Abstract
Recent vision language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage”, where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to the over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address these issues, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning, as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-Unlearn. WARNING: There exist AI generations that may be offensive in nature.
@@ 1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023a). A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023, 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a). While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023). Specifically, images can function as a “foreign language” (Pi et al., 2024), creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025). The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024). One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024a), which substantially improves robustness against unsafe queries and jailbreaking attacks. The surprising effectiveness of safety SFT has sparked growing interest in re-evaluating its reliability. However, a key limitation of such safety fine-tuning is the issue of over-prudence (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), where VLMs become overly conservative and unnecessarily reject benign queries.
The over-prudence may indicate that these models are overly safe, avoiding responses even to benign queries. Alternatively, the observed safety may be illusory, as current fine-tuning fails to guarantee reliability, creating a false sense of safety. Thus, we ask:
@@
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed fine-tuned on VLGuard (Zong et al., 2024)) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in an unsafe query) bypasses the safety mechanism learned from fine-tuning on VLGuard, even though the model correctly rejects the original query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations from fine-tuning dataset biases: Certain words are disproportionately linked with specific safety labels, such as “Share” strongly correlated with rejection, while “What” highly associated with non-rejection. (d) Effectiveness of unlearning-based safety fine-tuning: Using unlearning methods NPO (Zhang et al., 2024b) and RMU (Li et al., 2024a) to do the safety fine-tune on the VLM, which could enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
@@
In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs. We uncover a “safety mirage” in VLMs safety fine-tuning, where the seemingly robust performance largely stems from spurious correlations between specific words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset. If an adversary identifies these spurious correlations, a simple one-word modification–which we refer to as the “one-word attack”–can effectively jailbreak safety fine-tuned VLMs, enabling them to regenerate unsafe content. Additionally, the input-rejection shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs. Similar to the one-word attack, a one-word modification at test time can easily trigger this shortcut, causing the model to overgeneralize rejection and unnecessarily refuse benign queries. In Fig. 1(a)-(c), we provide a schematic overview illustrating: (a) The proposed one-word attack on a safety fine-tuned VLMs LLaVA-v1.5-7B-Mixed (Zong et al., 2024); (b) The amplified over-prudence issue by one-word modification; (c) The identified spurious correlations in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
@@
Building on our identification of spurious correlations in VLMs safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU preserves essential knowledge while avoiding unintended disruptions to causally unrelated information (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021). We propose adapting MU to VLMs safety fine-tuning as a more robust alternative to traditional supervised approaches. Instead of enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLMs safety fine-tuning in prior work (Chakraborty et al., 2024; Chen et al., 2025; Huo et al., 2025), its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024b) and RMU (representation misdirection for unlearning) (Li et al., 2024a), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
@@
①We revisit the problem of safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases, specifically, spurious correlations between textual questions and safety labels in the fine-tuning dataset.
②From the attack perspective, we show that safety fine-tuned VLMs remain vulnerable to jailbreaking when adversaries exploit these spurious correlations. We propose a simple yet effective one-word attack, which substitutes highly frequent query words associated with rejection responses with those linked to normal model outputs. Furthermore, we demonstrate that spurious correlations also underlie the over-prudence phenomenon, causing models to unnecessarily reject benign inputs.
③From the defense perspective, we show that machine unlearning (MU) offers a promising solution to mitigate the effects of spurious correlations in fine-tuning data. The key idea is that MU removes the influence of unsafe responses without relying on spurious feature–label shortcuts.
④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024a), MM-SafetyBench (Liu et al., 2024c), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
@@ 2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs (Liu et al., 2023, 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023a; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024b; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023). Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024c). Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023b; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms (Wang et al., 2024b; Chen et al., 2023b; Pi et al., 2024; Gou et al., 2024; Ding et al., 2024). The latter focuses on red-teaming dataset curation (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024; Zong et al., 2024; Li et al., 2024c), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
Machine unlearning in VLMs. MU (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024b; Li et al., 2024a; Yao et al., 2024). In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms (Liu et al., 2024d; Dontsov et al., 2024; Ma et al., 2024). For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation (Chen et al., 2025; Huo et al., 2025; Chakraborty et al., 2024). Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
@@ 3 Preliminary and Problem Statement
Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024c) highlight the importance of safety alignment in VLMs across both textual and visual modalities. Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs. For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign. SPA-VL (Zhang et al., 2024a) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response. Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023b; Röttger et al., 2023). In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
We begin by formulating the problem of VLM safety fine-tuning. Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard and SPA-VL). Let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, consisting of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning. For a VLM parameterized by $\bm{\theta}$}, the safety fine-tuning problem can be formulated as:
@@
where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and $\gamma\geq 0$ is a regularization parameter that balances safety alignment with general performance preservation.
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc (Zong et al., 2024; Taori et al., 2023)). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. However, asking the initial question by starting with the word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger refusals even for benign queries.
Safety mirage: Motivation and problem of interest. Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024a). The seemingly ‘robust’ safety observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement, Fig. 2 & A1 provides motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. For example, for the VLGuard fine-tuned models, unsafe queries prefixed with the innocuous question word “What” successfully bypass the safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the question word is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
Examples in Figs. 1, 2, and A1 suggest that fine-tuning VLMs on safety datasets creates a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries. Thus, our work focuses on the following key research questions: (a) What is the root cause of the “safety mirage” in VLM safety fine-tuning? (b) What can be improved to mitigate the “safety mirage”? We explore these questions in Sec. 4 and Sec. 5, respectively.